Prove that

lim n→∞ [ 1/ (2n-1) + 1/ (4n-22) + 1/ (6n-32) +.... + 1/n ] = π /2

1
Expert's answer
2021-06-07T17:23:58-0400

"lim _{n→∞} [ \frac{1}{\sqrt{2n-1}}+\frac{1}{\sqrt{4n-2^2}}+\frac{1}{\sqrt{6n-3^2}}+•••+\frac{1}{n}]=lim _{n→∞}\sum_{k=1 } ^n [ \frac{1}{\sqrt{2kn-k^2}}]\\
=lim _{n→∞}\sum_{k=1 } ^n \frac{1}{n}
 \frac{1}{\sqrt{2(\frac{k}{n})-(\frac{k}{n})
^2}}\\
\text{By integral test,}\\
\int_0^1 \frac{1}{\sqrt{2x-x^2}}dx\\
=\int_0^1 \frac{1}{\sqrt{1-(x-1)^2}}dx\\
Put \space x-1=t\implies dx=dt\\
\int_{-1}^0 \frac{1}{\sqrt{1-t^2}}dt\\
=sin^{-1}t]_{-1}^0\\
=sin^{-1}(0)-sin^{-1}(-1)\\
=0-(-\frac{\pi}{2})\\
=\frac{\pi}{2}"